OpenStudy (anonymous):

Use the distributive property to simplify the expression below: The quantity of negative 5 times x plus 20 all divided by negative 5 −x − 4 x + 20 x + 4 x − 4

OpenStudy (nincompoop):

this is an example of distributive property a(b+c) = ab + bc another example would be 3 (ab - 3cd) = 3ab -9cd what essentially is happening here is that whatever is outside the parenthesis is being distributed to whatever is inside the parenthesis by means of multiplication. another example using arithmetic example would be 3 (5 + 2) = 15 + 6 = 21 see we can also do this 3 (5 + 2) = 3 (7) = 21 they yielded the same response, I just showed you how distributive property works. It doesn't give you any different than if you followed the ordered of operation when applying to arithmetic figures.